Daisy is pulling a 20 kg sled to the right over a rough surface with a single horizontal rope. The tension in the rope is 100.8 N and the force of friction between the surface and the sled is 65.7 N . What is the acceleration of the sled? Express your answer in meters per second squared and round to one decimal place.
Answer #1

F=Net Horizontal Force

T=Tension in the rope

f=Friction force

m=Mass of the sled

F= T-f= 100.8-65.7= 35.1 N

According to Newton's 2 law

F=ma [Where a is acceleration the body]
a= F/m= 35.1/20= 1.8 m/s2